### v3

@@ Abstract
In Bayesian quantile regression, the asymmetric Laplace (AL) distribution is the most commonly used likelihood. The reason for this choice is not its plausibility as a data-generating distribution but that the corresponding maximum likelihood estimator is identical to the classical estimator by Koenker and Bassett (1978). In that sense, the AL likelihood is just a working likelihood. While point estimation is consistent, credible intervals tend to have poor frequentist coverage unless the data-generating distribution happens to be the AL distribution. Yang et al. (2016) proposed an adjustment to the posterior covariance matrix that produces asymptotically valid intervals. However, we show that this adjustment is sensitive to the AL distribution’s scale parameter and can lead to poor coverage when sample sizes are moderate. We propose using Infinitesimal Jackknife (IJ) standard errors (Giordano and Broderick, 2024), which require no resampling and can be obtained from a single MCMC run. We also propose a version of IJ standard errors for clustered data. Simulations and applications to real data show that IJ standard errors have good frequentist properties for both independent and clustered data. We provide an R package, IJSE, that computes IJ standard errors after estimation with the brms wrapper for Stan.
@@
VALID STANDARD ERRORS FOR BAYESIAN QUANTILE REGRESSION WITH CLUSTERED AND INDEPENDENT DATA
Abstract
@@ 1 Introduction
In frequentist quantile regression, point estimates can be obtained by minimizing an appropriate loss function through linear programming algorithms (Koenker & Bassett, 1978). Unlike this classic approach, Bayesian quantile regression requires a likelihood. There are several proposals including nonparametric likelihoods (e.g., Kottas & Krnjajić, 2009; Reich et al., 2010) and empirical likelihoods (e.g., Otsu, 2008; Yang & He, 2012). However, by far the most commonly applied likelihood is the asymmetric Laplace (AL) likelihood first employed by Yu and Moyeed (2001). The motivation for using an AL likelihood is that maximizing this likelihood corresponds to minimizing the loss function. The general approach of constructing a posterior through a loss function is discussed by, for instance, Syring and Martin (2019).
@@ 3 Infinitesimal Jackknife (IJ) standard errors
We propose using Infinitesimal Jackknife (IJ) standard errors introduced by Giordano and Broderick (2024), based on their earlier work (Giordano et al., 2018; Giordano, 2020; Giordano et al., 2020). Giordano (2020) motivates the IJ variance as an approximation to a resampling variance, such as bootstrap or jackknife. Since it would be computationally infeasible to perform MCMC estimation in a large number of resampled datasets, Giordano et al. (2020) propose using a linear approximation for the posterior means of the parameters in resampled data. Let $w$ be an $n$-dimensional weight vector, with element $w_{i}$ equal to the number of times unit $i$ is represented in the resampled data. The part of the posterior distribution that changes in resampled data is the likelihood function, and we can express the log-likelihood for a given weight vector as
@@
and the corresponding Infinitesimal Jackknife (IJ) variance estimator is
@@ 7 Concluding Remarks
We propose using Infinitesimal Jackknife (IJ) standard errors (Giordano & Broderick, 2024) because they have low relative errors regardless of the value that $\sigma$ has been set to. Coverage is also good except at extreme quantiles when $\sigma$ is such that the point estimates are biased. We have seen little discussion in the literature of bias due to a poor choice of $\sigma$. We showed that the bias tends to increase with $\sigma$, as the posterior distributions of the regression parameters become increasingly skewed, particularly for the intercept parameter.

### v4

@@ Abstract
Bayesian quantile regression typically uses the asymmetric Laplace (AL) distribution as working likelihood, not because it is a plausible data-generating distribution but because the corresponding maximum likelihood estimator is identical to the classical estimator by Koenker and Bassett (1978). While point estimation is consistent, credible intervals tend to have poor frequentist coverage. We propose using Infinitesimal Jackknife (IJ) standard errors (Giordano and Broderick, 2024), which require no resampling and can be obtained from a single MCMC run. Simulations and applications to real data show that IJ standard errors have good frequentist properties for both independent and clustered data. We provide an R package, IJSE, that computes IJ standard errors after estimation with the brms wrapper for Stan.
@@ 1 Introduction
Quantile regression is popular across a wide range of fields, including education and psychology (e,g., Betebenner, 2009; Castellano & Ho, 2013; Furno, 2011; Konstantopoulos et al., 2019; Zhu et al., 2024). In frequentist quantile regression, point estimates can be obtained by minimizing an appropriate loss function through linear programming algorithms (Koenker & Bassett, 1978). Unlike this classic approach, Bayesian quantile regression requires a likelihood. There are several proposals including nonparametric likelihoods (e.g., Kottas & Krnjajić, 2009; Reich et al., 2010) and empirical likelihoods (e.g., Otsu, 2008; Yang & He, 2012). However, by far the most commonly applied likelihood is the asymmetric Laplace (AL) likelihood first employed by Yu and Moyeed (2001). The motivation for using an AL likelihood is that maximizing this likelihood corresponds to minimizing the loss function. The general approach of constructing a posterior through a loss function is discussed by, for instance, Syring and Martin (2019).
@@ 3 Infinitesimal Jackknife (IJ) standard errors
Infinitesimal jackknife standard errors were originally proposed for frequentist estimators (Jaeckel, 1972) and applied in covariance structure analysis by Jennrich (2008). A Bayesian version, which we will simply denote IJ for brevity, was introduced by Giordano and Broderick (2024), based on their earlier work (Giordano et al., 2018; Giordano, 2020; Giordano et al., 2020). Giordano (2020) motivates the IJ variance as an approximation to a resampling variance, such as bootstrap or jackknife. Since it would be computationally infeasible to perform MCMC estimation in a large number of resampled datasets, Giordano et al. (2020) propose using a linear approximation for the posterior means of the parameters in resampled data. Let $w$ be an $n$-dimensional weight vector, with element $w_{i}$ equal to the number of times unit $i$ is represented in the resampled data. The part of the posterior distribution that changes in resampled data is the likelihood function, and we can express the log-likelihood for a given weight vector as
@@
and the corresponding infinitesimal jackknife variance estimator is
@@ 7 Concluding Remarks
We propose using Bayesian Infinitesimal Jackknife (IJ) standard errors (Giordano & Broderick, 2024) because they have low relative errors regardless of the value that $\sigma$ has been set to. Coverage is also good except at extreme quantiles when $\sigma$ is such that the point estimates are biased. We have seen little discussion in the literature of bias due to a poor choice of $\sigma$. We showed that the bias tends to increase with $\sigma$, as the posterior distributions of the regression parameters become increasingly skewed, particularly for the intercept parameter.